# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ V-B1 Dynamic Scattering Arrays
We illustrate a numerical example in which (40) is optimized to realize a MIMO EM precoder. In particular, we consider a MIMO communication in which the DSA is adopted at the transmitter as illustrated in Fig. 8. A receiving user equipped with a standard ULA with $K=20$ elements spaced of $\lambda/2$ is located at 101010\,10m in non-line-of-sight (NLOS) condition. The simulated NLOS channel consists of 5 multi-paths caused by the reflection of 5 scatterers in the environment located at angles $\phi=\{-43^{\circ},-14^{\circ},14^{\circ},43^{\circ},72^{\circ}\}$ and distance 555\,5m. The corresponding strongest singular values of the channel are $(-35.6{\,\text{dB}},-42.4{\,\text{dB}},-43.7{\,\text{dB}},-46.3{\,\text{dB}},-82{\,\text{dB}},\ldots)\,$, which has clearly rank $r=4$. The DSA is equipped with $N_{\text{S}}=121$ Hertzian reconfigurable scattering dipoles deployed in 5 concentric rings and spaced apart of $\lambda/4$. The carrier frequency is 282828\,28GHz so that the total size of the DSA is 3.23.23.2\,3.2cm. It is well-known from MIMO theory that for a given generic propagation scenario characterized by rank $r$, up to $r$ parallel orthogonal links (layers) can be established between the transmitter and the receiver on which independent data stream can be transmitted [42]. To exploit all of them, the number of active antennas (i.e., RF chains) must be $N_{\text{A}}=r$ and the DSA must implement a suitable precoding strategy, i.e., act as an EM precoder. Specifically, setting the target end-to-end channel to be diagonal and matching the singular values of the link ${\bf{H}_{\text{o}}}=\alpha\mathbf{G}_{\text{R}}$ (channel diagonalization), must be
$\left(\jmath\,{\rm diag}\left(\bm{\theta}\right)-{\bf{Z}}\right)^{-1}\propto{\bf{U}}^{{\mathrm{H}}}$ (41)
@@
where ${\bf{U}}^{{\mathrm{H}}}$ represents the optimal precoding matrix. The resulting end-to-end matrix $\hat{\mathbf{H}}_{\text{o}}$, obtained by solving (41) numerically while maximizing the radiated power $P_{\text{T}}$ (see [30] for the details), is reported in Fig. 8. The comparison between the diagonal and off-diagonal values indicates that the coupling between different layers is completely negligible, i.e., the channel is almost perfectly diagonalized. Compared to the actual singular values of the channel, the intensity of the diagonal elements of $\hat{\mathbf{H}}_{\mathrm{o}}$ exhibit the same behavior with a loss of about 666\,6dB that can be ascribed to the fact that the antenna is not ideal. The designed DSA implements the optimal precoding at the EM level by using no more than $N_{\text{A}}=r$ RF chains (i.e., the minimum possible value) which is much simpler and less energy-consuming than any conventional full digital or hybrid solution by considering that the reconfigurable scattering elements are passive. Further examples of DSA design for the realization of superdirective beamforming and multi-user multiple-input single-output (MISO) can be found in [30]. The main challenge from the SP perspective lies in finding efficient techniques to tackle high-dimensional optimization problems constrained by EM laws like that in (40) for a large number of reconfigurable scattering elements.
@@ V-B2 Stacked Intelligent Metasurfaces
Recently, a new technology known as SIM has gained significant attention within the research community due to its advantages and potential applications [33]. A SIM consists of a sealed vacuum structure that incorporates multiple layers of metasurfaces, each containing many reconfigurable meta-atoms, i.e., small reconfigurable scattering elements, which are interconnected with a controller, e.g., a field programmable gate array (FPGA). Operating on the well-established Huygens-Fresnel principle, when an EM wave interacts with a meta-atom in any given layer, it acts as a secondary point source, illuminating all subsequent meta-atoms in the following layer. This architectural approach, which can be seen as a particular case of DSAs, with cells organized in layers and coupling occurring in only one direction, provides greater flexibility in EM waveform manipulation and enhanced spatial-domain gain when compared to conventional MIMO architectures, e.g., transmitarrays and hybrid MIMO antennas [47]. It significantly reduces the number of required RF chains, hence allowing for improved EM-level processing with very limited energy consumption [48].
@@
Interestingly, thanks to their peculiar layered structure, several optimization algorithms for determining the optimal SIM transmission coefficients have been inspired by the well-known error backpropagation algorithm used for DNN training, hence leveraging their particular hardware architecture [49]. These algorithms allow for the optimization of the meta-atom configuration and the phase shifts imposed on EM waves passing through these structures. Such optimizations enable the implementation of various wave-domain functionalities, such as beamforming, precoding, and combining. For instance, [49] applied deep reinforcement learning to jointly optimize the phase shifts of SIMs and the transmit power allocation. In a complementary approach, [50] proposed a projected gradient descent (PGD) method to iteratively optimize the phase shifts of SIMs at the transceivers layer-by-layer, aimed at minimizing the channel cutoff rate under practical modulation constraints. Furthermore, [51] introduced a hybrid digital and wave-domain SIM-based MIMO transceiver architecture and proposed a projected gradient ascent (PGA) method that jointly optimizes the phase shifts of the SIM, alongside digital transmit precoding and receive combining. Further works as [52, 53] employed backpropagation-inspired approaches to find the optimal SIMs configurations for specific tasks like channel estimation, image reconstruction, or DoA estimation.
Figure 9: SIM-based antenna architecture, comprising a ULA, $L$ metasurface layers, and a FPGA for reconfigurability. Example of performance comparison of the SIM-based DoA estimator and the conventional approach via a 2D digital DFT [53].
@@
To illustrate, let us consider a SIM-based receiving antenna as depicted in Fig. 9, where a receiving SIM composed of $L$ metasurface layers is followed by a conventional antenna array, such as an ULA, consisting of $N_{\text{A}}$ active antennas, being $N_{\text{A}}$ also the number of receiving RF chains. For simplicity, we assume that every metasurface layer composing the SIM is equipped with the same number $M$ of meta-atoms, where $M\geq N_{\text{A}}$. Moreover, we assume that each layer of the SIM has an isomorphic square lattice arrangement and we model each metasurface layer as a uniform planar array [49]. Specifically, the distance between adjacent meta-atoms is denoted as $d_{\mathrm{atom}}$, the area of each meta-atom is $A$, and the spacing between each metasurface layer is represented by $d_{\mathrm{layer}}$.
@@
The transmission coefficient of the $m$-th meta-atom in the $l$-th metasurface layer is represented by $\phi_{m}^{(l)}=e^{\jmath\theta_{m}^{(l)}}$, with $\theta_{m}^{(l)}\in[0,2\pi)$, for $m=1,2,\dots,M$ and $l=1,2,\dots,L$. Consequently, the matrix of transmission coefficients for the $l$-th metasurface layer can be expressed as $\bm{\Phi}^{(l)}={\rm diag}\left(\phi_{1}^{(l)},\phi_{2}^{(l)},\ldots,\phi_{M}^{(l)}\right)\in\mathbb{C}^{M\times M}$. Moreover, we denote the intra-layer propagation matrix of the SIM, which describes the transmission coefficients from the $(l-1)$-th transmit metasurface layer to the $l$-th layer, as $\mathbf{W}^{(l)}\in\mathbb{C}^{M\times M}$ for $l=1,2,\dots,L$. According to Rayleigh-Sommerfeld diffraction theory [35], the $(m,i)$-th entry of $\mathbf{W}^{(l)}$ is expressed as
@@
$w_{m,i}^{(l)}=\frac{A\,d_{\mathrm{layer}}}{d_{m,i}^{(l)}}\left(\frac{1}{2\pi d_{m,i}^{(l)}}-\frac{\jmath}{\lambda}\right)e^{\jmath k_{0}\,d_{m,i}^{(l)}},\quad l=1,2,\dots,L$ (42)
@@
where $d_{m,i}^{(l)}$ is the transmission distance from the $i$-th meta-atom of the $(l-1)$-th metasurface layer to the $m$-th meta-atom of the $l$-th layer. Therefore, the overall effect of the receiving SIM can be represented as
$\mathbf{R}(\bm{\theta})=\bm{\Phi}^{(L)}\mathbf{W}^{(L)}\ldots\bm{\Phi}^{(2)}\mathbf{W}^{(2)}\bm{\Phi}^{(1)}\mathbf{W}^{(1)}\in\mathbb{C}^{M\times M}$
@@
where $\bm{\theta}=\left\{\theta_{m}^{(l)}\right\}$, for $m=1,2,\ldots,M$, $l=1,2,\ldots,L$, collects all the parameters to be optimized. In addition, we assume that the receiving ULA is positioned close to the last SIM layer, that is, within its reactive near-field propagation region. Consequently, the wireless interaction between the receive ULA and the last layer of the SIM is characterized by the matrix ${\bf{H}}_{\text{R}}=\left\{h_{m,s}\right\}\in\mathbb{C}^{N_{\text{A}}\times M}$, whose elements can be derived by substituting $d_{m,i}^{(l)}$ in (42) with $r_{m,s}^{(1)}$, i.e., the distance between the $m$-th meta-atom of the last layer of the SIM and the generic $s$-th receiving antenna. Note that, despite matrices $\bm{\Phi}^{(l)}$ are diagonal, the coupling effect at each layer makes the overall reflection matrix ${\bf{R}}(\bm{\theta})$ non-diagonal thus increasing the variables available for the optimization. When optimizing the phase profiles of the SIM for a given task, e.g., to realize an EM MIMO precoder, perform localization or channel estimation, the optimization problem that needs to be addressed is often a particular case of (40), which typically is in the form [33]
@@
$\underset{\bm{\theta}}{\operatorname{minimize}}$ $\quad\bm{\Gamma}=\left\|\,\mathbf{R}(\bm{\theta})\,{\bf{H}}_{\text{R}}-{\bf{H}_{\text{o}}}\right\|_{\mathrm{F}}^{2}$ (43)
@@
where ${\bf{H}_{\text{o}}}$ is a target response matrix that depends on the specific functionality to be achieved. Thanks to the layered structure of the SIM, the solution to (43) can be found thanks to a gradient descent algorithm based on error back-propagation, as outlined in Algorithm 1.
1:  Input: $\mathbf{W}^{(l)},\,l=1,\dots,L,\,{\bf{H}_{\text{o}}}$
@@
For instance, an interesting application of the algorithm mentioned above can be found in [53], where the Authors formulate an optimization problem in the form of (43) to compute the 2D DFT in the wave domain, hence propose a gradient descent algorithm to obtain a near-optimal solution for the SIM transmission coefficients, minimizing the Frobenius norm of the approximation error between the SIM’s end-to-end transfer function and the ideal 2D DFT matrix ${\bf{H}_{\text{o}}}$. The final objective of [53] is to leverage the 2D DFT, resulting from the propagation of the incident wave through the optimized SIM, to estimate the DoA of the signal at the receiving antenna array placed after the SIM. In the absence of noise, the EM waves propagating through the SIM are automatically focused on the specific receiving antenna corresponding to the on-grid DoA estimate of the incoming signal thanks to the 2D DFT operation. As a result, the DoA of the incoming signal can be readily estimated by simply measuring the energy distribution across the receiver antenna array.
In this regard, in Fig. 9-right the performance of such a DoA estimator is reported from [53], in comparison to the conventional method based on digital beamforming. Specifically, the mean squared error (MSE) of the electrical DoA angles $\psi_{\text{x}}$ and $\psi_{\text{y}}$ in the x- and y-directions is reported. The results show that the SIM-based estimator, operating with $N_{\text{A}}=4\times 4=16$ antennas and collecting the received signals over $T$ = $64$ snapshots before DoA estimation, performs comparably to the digital beamforming-based method with the same number of receiver antennas at large signal-to-noise ratios. Additionally, Fig. 9-right plots the MSE of a $32\times 32$ digital receiver array with a single snapshot, which provides a lower bound for the SIM-based scheme.
The results indicate that the SIM-based estimator performs comparably to the digital method, with the benefit of relying on energy detection and wave-based signal processing instead of coherent receivers and complex digital processing. Moreover, despite the superior accuracy, the reference digital approach demands a significantly larger array aperture with much more RF chains. In contrast, the SIM-based estimator achieves effective DoA estimation with a smaller array, demonstrating its practical advantage in resource-constrained scenarios. Details regarding the settings used to obtain the numerical results presented in Fig. 9 are available in [53].
@@
Another interesting application can be found in [32], which introduces a novel microwave-based diffractive deep neural network ($\mathrm{D}^{2}\mathrm{NN}$) designed to classify handwritten metallic digits. The proposed architecture consists of a SIM comprising $L=3$ sequential metasurface layers, each of them having $M=1024$ meta-atoms capable of introducing phase modulation to the impinging EM field. Input data, represented as engraved metal digit patterns, are illuminated by EM waves, propagating through the $\mathrm{D}^{2}\mathrm{NN}$ layers, and focusing energy at specific points on the target plane, hence enabling classification. The phase responses of each layer are designed using a stochastic gradient descent algorithm to minimize a cross-entropy loss function based on a training dataset. The system workflow is depicted in Fig. 10-left, which shows the overall SIM structure, including the input plane of engraved digits and the output layer with the EM field distribution which focuses the energy onto designated points at the target plane for direct classification. The optimized phase distributions of the three diffractive layers, presented in Fig. 10-middle, illustrate how the SIM processes input data by modulating the phase of the transmitted EM waves. These distributions reveal that the $\mathrm{D}^{2}\mathrm{NN}$ architecture progressively refine the wavefront, enabling precise focusing at the output plane and effectively mapping the input patterns into distinct classification categories. The phase patterns also highlight the network’s spatial symmetry and its ability to represent complex mappings within its design constraints. Finally, Fig. 10-right examines the classification accuracy as a function of the relative scale of the intermediate diffractive layers, i.e., based on the number of meta-atoms per layer side. The results indicate that the accuracy improves with increasing scale, plateauing at approximately $90.4\%$ for a three-layer configuration, validating the efficiency of the gradient descent algorithm-based optimization in achieving high classification accuracy using a SIM.

### v3

@@ V-B1 Dynamic Scattering Arrays
Hereafter, we illustrate a numerical example in which (42) is optimized to realize a MIMO EM precoder. In particular, we consider a MIMO communication in which the DSA is adopted at the transmitter as illustrated in Fig. 8. A receiving user equipped with a standard ULA consisting of $K=20$ elements spaced at $\lambda/2$ is located at a distance of 10 m under non-line-of-sight (NLOS) conditions. The simulated NLOS channel consists of 5 multi-paths caused by the reflection of 5 scatterers in the environment corresponding to the following strongest singular values of the channel $(-35.6\,\mathrm{dB},-42.4\,\mathrm{dB},-43.7\,\mathrm{dB},-46.3\,\mathrm{dB},-82\,\mathrm{dB},\ldots)\,$, which has clearly rank $r=4$. The DSA is equipped with $N_{\text{S}}=121$ Hertzian reconfigurable scattering dipoles deployed in 5 concentric rings and spaced apart of $\lambda/4$. The carrier frequency is $28\,\mathrm{GHz}$ so that the total size of the DSA is $3.2\,\mathrm{cm}$. It is well known from MIMO theory that, in a generic propagation scenario characterized by rank $r$, up to $r$ parallel orthogonal links (or layers) can be established between the transmitter and the receiver, enabling the transmission of $r$ independent data streams [41]. To exploit all of them, the number of active antennas (i.e., RF chains) must be $N_{\mathrm{A}}=r$ and the DSA must implement a suitable precoding strategy, i.e., act as an EM precoder. Specifically, setting the target end-to-end channel to be diagonal and matching the singular values of the link ${\mathbf{H}_{\text{o}}}=\alpha\,\mathbf{G}_{\text{R}}{\mathbf{Q}}$ (channel diagonalization), it must be
$\left(\mathbf{Z}_{\text{L}}(\bm{\theta})+{\mathbf{Z}}\right)^{-1}\propto{\mathbf{W}}\,$, (43)
where ${\mathbf{W}}={\mathbf{U}}^{-1}$ represents the optimal precoding matrix. The approximating end-to-end channel matrix $\hat{\mathbf{H}}_{\text{o}}$, obtained by solving (42) numerically while maximizing the radiated power $P_{\text{T}}$ (see [26] for the details), is reported in Fig. 8.
@@
The comparison between the diagonal and off-diagonal values indicates that the coupling between different layers is completely negligible, i.e., the channel is almost perfectly diagonalized. Compared to the actual singular values of the channel, the intensity of the diagonal elements of $\hat{\mathbf{H}}_{\mathrm{o}}$ exhibit the same behavior with a loss of about 66\,dB that can be ascribed to the fact that the antenna is not ideal. The designed DSA implements optimal precoding at the EM level using no more than $N_{\mathrm{A}}=r$ RF chains—the minimum possible—making it significantly simpler and more energy-efficient than conventional full-digital or hybrid solutions, especially considering that the reconfigurable scattering elements are passive. Further examples of DSA design for the realization of superdirective beamforming and multi-user multiple-input single-output (MISO) can be found in [26]. The main challenge from the SP perspective lies in finding efficient techniques to tackle high-dimensional optimization problems constrained by EM laws like that in (42) for a large number of reconfigurable scattering elements.
@@ V-B2 Stacked Intelligent Metasurfaces
Recently, a new technology known as SIM has gained significant attention within the research community due to its advantages and potential applications [29]. A SIM consists of a sealed vacuum structure that incorporates multiple layers of metasurfaces, each containing many reconfigurable meta-atoms, i.e., small reconfigurable scattering elements, which are interconnected with a controller (e.g., an FPGA) whose purpose is to implement the optimization algorithm (see the example below) and change the response of each meta-atom, typically consisting in a two-state phase shift. Operating on the well-established Huygens-Fresnel principle, when an EM wave interacts with a meta-atom in any given layer, it acts as a secondary point source, illuminating all subsequent meta-atoms in the following layer. This architectural approach, which can be seen as a particular case of DSAs, with cells organized in layers and coupling occurring in only one direction, provides greater flexibility in EM waveform manipulation and enhanced spatial-domain gain when compared to conventional MIMO architectures, e.g., transmitarrays and hybrid MIMO antennas [47]. It significantly reduces the number of required RF chains, hence allowing for improved EM-level processing with very limited energy consumption [29].
@@
Interestingly, thanks to their peculiar layered structure, several optimization algorithms for determining the optimal SIM transmission coefficients have been inspired by the well-known error backpropagation algorithm used for DNN training, hence leveraging their particular hardware architecture [48]. These algorithms allow for the optimization of the meta-atom configuration and the phase shifts imposed on EM waves passing through these structures. Such optimizations enable the implementation of various wave-domain functionalities, such as beamforming, precoding, and combining. For instance, [48] applied deep reinforcement learning to jointly optimize the phase shifts of SIMs and the transmit power allocation. In a complementary approach, [49] proposed a projected gradient descent (PGD) method to iteratively optimize the phase shifts of SIMs at the transceivers layer-by-layer, aimed at minimizing the channel cutoff rate under practical modulation constraints. Furthermore, [50] introduced a hybrid digital and wave-domain SIM-based MIMO transceiver architecture and proposed a projected gradient ascent (PGA) method that jointly optimizes the phase shifts of the SIM, alongside digital transmit precoding and receive combining. Further works, such as [51, 52], employed backpropagation-inspired approaches to determine the optimal SIM configurations for specific tasks, including channel estimation, image reconstruction, and DoA estimation.
@@
To illustrate, consider a SIM-based receiving antenna as shown in Fig. 9, where a receiving SIM composed of $L$ metasurface layers is followed by a conventional antenna array, such as a ULA, consisting of $N_{\mathrm{A}}$ active antennas, with $N_{\mathrm{A}}$ also representing the number of receiving RF chains. For simplicity, we assume that every metasurface layer composing the SIM is equipped with the same number $M$ of meta-atoms, where $M>N_{\mathrm{A}}$. [^9] The condition $M>N_{\mathrm{A}}$ is not strictly necessary; however, it is generally desirable. In practice, $N_{\mathrm{A}}$ is kept small to limit the number of RF chains, thereby reducing hardware complexity and power consumption, whereas $M$ is chosen to be large to increase the DoF available for processing the EM wave. Moreover, we assume that each layer of the SIM has an isomorphic square lattice arrangement and we model each metasurface layer as a uniform planar array [48]. Specifically, the distance between adjacent meta-atoms is denoted as $d_{\mathrm{atom}}$, the area of each meta-atom is $A$, and the spacing between each metasurface layer is represented by $d_{\mathrm{layer}}$.
Figure 9: SIM-based antenna architecture, comprising a ULA, $L$ metasurface layers, and a controller for reconfigurability. Example of performance comparison of the SIM-based DoA estimator and the conventional approach via a 2D digital DFT [52].
@@
The transmission coefficient of the $m$-th meta-atom in the $l$-th metasurface layer is represented by $\phi_{m}^{(l)}=e^{\jmath\theta_{m}^{(l)}}$, with $\theta_{m}^{(l)}\in[0,2\pi)$, for $m=1,2,\dots,M$ and $l=1,2,\dots,L$. Consequently, the matrix of transmission coefficients for the $l$-th metasurface layer can be expressed as $\bm{\Phi}^{(l)}={\rm diag}\left(\phi_{1}^{(l)},\phi_{2}^{(l)},\ldots,\phi_{M}^{(l)}\right)\in\mathbb{C}^{M\times M}$. Moreover, we denote the intra-layer propagation matrix of the SIM, which describes the transmission coefficients from the (l1)(l-1)-th transmit metasurface layer to the $l$-th layer, as $\mathbf{W}^{(l)}\in\mathbb{C}^{M\times M}$ for $l=1,2,\dots,L$. According to Rayleigh-Sommerfeld diffraction theory [35], the $(m,i)$-th entry of $\mathbf{W}^{(l)}$ is expressed as
@@
$w_{m,i}^{(l)}=\frac{A\,d_{\mathrm{layer}}}{d_{m,i}^{(l)}}\left(\frac{1}{2\pi d_{m,i}^{(l)}}-\frac{\jmath}{\lambda}\right)e^{\jmath k_{0}\,d_{m,i}^{(l)}},\quad l=1,2,\dots,L\,$, (44)
@@
where $d_{m,i}^{(l)}$ is the transmission distance from the $i$-th meta-atom of the (l1)(l-1)-th metasurface layer to the $m$-th meta-atom of the $l$-th layer. Therefore, the overall effect of the receiving SIM can be represented as
$\mathbf{R}(\bm{\theta})=\bm{\Phi}^{(L)}\mathbf{W}^{(L)}\ldots\bm{\Phi}^{(2)}\mathbf{W}^{(2)}\bm{\Phi}^{(1)}\in\mathbb{C}^{M\times M}\,$, (45)
@@
where $\bm{\theta}=\left\{\theta_{m}^{(l)}\right\}$, for $m=1,2,\ldots,M$, $l=1,2,\ldots,L$, collects all the parameters to be optimized. In addition, we assume that the receiving ULA is positioned close to the last SIM layer, that is, within its reactive near-field propagation region. Consequently, the wireless interaction between the receive ULA and the last layer of the SIM is characterized by the matrix ${\mathbf{H}}_{\text{R}}=\left\{h_{m,s}\right\}\in\mathbb{C}^{N_{\mathrm{A}}\times M}$, whose elements can be derived by substituting $d_{m,i}^{(l)}$ in (44) with $r_{m,s}^{(1)}$, i.e., the distance between the $m$-th meta-atom of the last layer of the SIM and the generic $s$-th receiving antenna. Note that, despite matrices $\bm{\Phi}^{(l)}$ being diagonal, the coupling effect at each layer makes the overall reflection matrix ${\mathbf{R}}(\bm{\theta})$ non-diagonal, thus increasing the variables available for the optimization. When optimizing the phase profiles of the SIM for a given task, e.g., to realize an EM MIMO precoder, perform localization or channel estimation, the optimization problem that needs to be addressed is often a particular case of (42), which typically is in the form [29]
@@
$\underset{\bm{\theta}}{\operatorname{minimize}}$ $\quad\bm{\Gamma}=\left\|{\mathbf{H}}_{\text{R}}\,\mathbf{R}(\bm{\theta})-{\mathbf{H}_{\text{o}}}\right\|_{\mathrm{F}}^{2}\,$, (46)
@@
where ${\mathbf{H}_{\text{o}}}\in\mathbb{C}^{N_{\mathrm{A}}\times M}$ is a target response matrix that depends on the specific functionality to be achieved. Thanks to the layered structure of the SIM, the solution to (46) can be found thanks to a gradient descent algorithm based on error back-propagation, as outlined in Algorithm 1.
1:Input: $\mathbf{W}^{(l)},\,l=1,\dots,L,\,{\mathbf{H}_{\text{o}}}$
@@
For instance, an interesting application of the algorithm mentioned above can be found in [52], where the Authors formulate an optimization problem in the form of (46) to compute the 2D DFT in the wave domain, hence propose a gradient descent algorithm to obtain a near-optimal solution for the SIM transmission coefficients, minimizing the Frobenius norm of the approximation error between the SIM’s end-to-end transfer function and the ideal 2D DFT matrix $\mathbf{H}_{\text{o}}$}. The final objective of [52] is to leverage the 2D DFT, resulting from the propagation of the incident wave through the optimized SIM, to estimate the DoA of the signal at the receiving antenna array placed after the SIM. In the absence of noise, the EM waves propagating through the SIM are automatically focused on the specific receiving antenna corresponding to the on-grid DoA estimate of the incoming signal thanks to the 2D DFT operation. As a result, the DoA of the incoming signal can be readily estimated by simply measuring the energy distribution across the receiver antenna array.
In this regard, in Fig. 9-right the performance of such a DoA estimator is reported from [52], in comparison to the conventional method based on digital beamforming. Specifically, the mean squared error (MSE) of the electrical DoA angles $\psi_{\text{x}}$ and $\psi_{\text{y}}$ in the x- and y-directions is reported. The results show that the SIM-based estimator, operating with $N_{\mathrm{A}}=4\times 4=16$ antennas and collecting the received signals over $T$ = $64$ snapshots before DoA estimation, performs comparably to the digital beamforming-based method with the same number of receiver antennas at large signal-to-noise ratios. Additionally, Fig. 9-right plots the MSE of a $32\times 32$ digital receiver array with a single snapshot, which provides a lower bound for the SIM-based scheme, since it directly resolves the same angular grid in a single snapshot that the SIM-based method synthesizes over $T=64$ snapshots with a $4\times 4$ array. The results indicate that the SIM-based estimator performs comparably to the digital method, with the benefit of relying on energy detection and wave-based signal processing instead of coherent receivers and complex digital processing. Moreover, despite the superior accuracy, the reference digital approach demands a significantly larger array aperture with many more RF chains. In contrast, the SIM-based estimator achieves effective DoA estimation with a smaller array, demonstrating its practical advantage in resource-constrained scenarios. Details regarding the settings used to obtain the numerical results presented in Fig. 9, including the use of different electrical angles for $\psi_{x}$ and $\psi_{y}$ estimation, resulting in slightly different MSE values, are available in [52].
@@
Another interesting application can be found in [28], which introduces a novel microwave-based diffractive deep neural network ($\mathrm{D}^{2}\mathrm{NN}$) designed to classify handwritten metallic digits. The proposed architecture consists of a SIM comprising $L=3$ sequential metasurface layers, each of them having $M=1024$ meta-atoms capable of introducing phase modulation to the impinging EM field. Input data, represented as engraved metal digit patterns, are illuminated by EM waves, propagating through the $\mathrm{D}^{2}\mathrm{NN}$ layers, and focusing energy at specific points on the target plane, hence enabling classification. The phase responses of each layer are designed using a stochastic gradient descent algorithm to minimize a cross-entropy loss function based on a training dataset. The system workflow is depicted in Fig. 10-left, which shows the overall SIM structure, including the input plane of engraved digits and the output layer with the EM field distribution which focuses the energy onto designated points at the target plane for direct classification. The optimized phase distributions of the three diffractive layers, presented in Fig. 10-middle, illustrate how the SIM processes input data by modulating the phase of the transmitted EM waves. These distributions reveal that the $\mathrm{D}^{2}\mathrm{NN}$ architecture progressively refines the wavefront, enabling precise focusing at the output plane and effectively mapping the input patterns into distinct classification categories.
Finally, Fig. 10-right examines the classification accuracy as a function of the relative scale of the intermediate diffractive layers, i.e., based on the number of meta-atoms per layer side. The results indicate that the accuracy improves with increasing scale, plateauing at approximately $90.4\%$ for a three-layer configuration, validating the efficiency of the gradient descent algorithm-based optimization in achieving high classification accuracy using a SIM.